OpenStudy (anonymous):

If y varies directly with x and y=7 and when x=56, then find. the constant of variation k

OpenStudy (diyadiya):

y varies directly with x means y = kx where k is proportionality constant

OpenStudy (diyadiya):

so here y=7 & x =56 substitute the value of y & x in the equation \[\Large y =kx\] can you find it ?

OpenStudy (diyadiya):

@rebecca1233 can you?

OpenStudy (anonymous):

so i times ?

OpenStudy (anonymous):

392?

OpenStudy (diyadiya):

\[\Large y=kx\]\[7= k \times 56\]to find k we divide both sides by 56 What do you get now?

OpenStudy (anonymous):

0.125

OpenStudy (diyadiya):

Right :)

OpenStudy (anonymous):

so thats the answer ?

OpenStudy (diyadiya):

Yeah 7/56=1/8=0.125

OpenStudy (anonymous):

thank you can you help me with a few more please

OpenStudy (diyadiya):

Will try :)

OpenStudy (anonymous):

If y varies directly with x and y=7 and when x=56, then find. the value of when x=4

OpenStudy (diyadiya):

So its the same question y varies directly with x so what will be the equation ?

OpenStudy (anonymous):

7/4?

OpenStudy (anonymous):

x=ky

OpenStudy (diyadiya):

Nope , y varies directly with x so y=kx if its x varies directly with y it will be x=ky understood ?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

y = kx 7=k*56 k=7/56 =1/8

OpenStudy (diyadiya):

For these type of questions ,First step is to find the value of k ok?

OpenStudy (anonymous):

ok

OpenStudy (diyadiya):

So now we have k =1/8 (or 0.125) consider the same equation y=k*x here x = 4 & k =0.125 so what is y = ?

OpenStudy (anonymous):

u divide right ?

OpenStudy (diyadiya):

y = k* x which all values you have ?

OpenStudy (diyadiya):

ok tell me what are we trying to find ..?

OpenStudy (anonymous):

idk it says the value of when x=4

OpenStudy (diyadiya):

you probably read the question wrong ,it will be ' the value of y when x=4"

OpenStudy (diyadiya):

so we are finding y here

OpenStudy (anonymous):

not on my paper it doesnt say y or anything

OpenStudy (diyadiya):

it might be printing mistake " the value of when x=4 " doesn't make any sense

OpenStudy (diyadiya):

do you have the answer ?

OpenStudy (anonymous):

hno

OpenStudy (anonymous):

how do you do this

OpenStudy (anonymous):

give the equation of the line

OpenStudy (anonymous):